3. If 70 mL of 0.01M HCL is added to 150 mL of 0.1M acetic acid, pH 5.0, what is the resulting pH? Identify the appropriate acid and conjugate base and determine their concentrations in the final solution. 


1
Expert's answer
2021-11-25T01:29:15-0500

70*0.01=0.7

150*0.1=15

70+150 — 5.0

0.7+15 — x

x=15.7*5/220

X=0.36